# Ada-NAV: Adaptive Trajectory Length-Based Sample Efficient Policy Learning for Robotic Navigation

### v4

@@ Abstract
Trajectory length stands as a crucial hyperparameter within reinforcement learning (RL) algorithms, significantly contributing to the sample inefficiency in robotics applications. Motivated by the pivotal role trajectory length plays in the training process, we introduce Ada-NAV, a novel adaptive trajectory length scheme designed to enhance the training sample efficiency of RL algorithms in robotic navigation tasks. Unlike traditional approaches that treat trajectory length as a fixed hyperparameter, we propose to dynamically adjust it based on the entropy of the underlying navigation policy. Interestingly, Ada-NAV can be applied to both existing on-policy and off-policy RL methods, which we demonstrate by empirically validating its efficacy on three popular RL methods: REINFORCE, Proximal Policy Optimization (PPO), and Soft Actor-Critic (SAC). We demonstrate through simulated and real-world robotic experiments that Ada-NAV outperforms conventional methods that employ constant or randomly sampled trajectory lengths. Specifically, for a fixed sample budget, Ada-NAV achieves an 18% increase in navigation success rate, a 20-38% reduction in navigation path length, and a 9.32% decrease in elevation costs. Furthermore, we showcase the versatility of Ada-NAV by integrating it with the Clearpath Husky robot, illustrating its applicability in complex outdoor environments.
@@ I Introduction
Autonomous robotic navigation is ubiquitous in tasks such as search-and-rescue [1], disaster relief[2], and planetary exploration applications [3]. While traditional navigation methods have been popular among the community [4, 5, 6], reinforcement learning (RL) [7] has been recently successful in robotic navigation tasks [8]. RL-based methods offer advantages in terms of being mapless, having lower dependence on the sensor accuracy, and possessing the ability to work with raw sensor data [8, 9, 10]. However, the benefits of RL come with their own set of intricate challenges, most prominently the issue of exploration in the environment during training. Traditional RL methods often hinge on the availability of a dense reward structure [11], which may rarely be available in complex, unpredictable environments [12, 13, 14, 15]. These challenges make the existing RL methods sample inefficient during training, especially for robotics applications where rewards are sparse [16, 17], since the agent does not receive constant feedback on its progression towards the goal.
@@
Figure 1: Robot navigation using policies trained with the REINFORCE policy gradient algorithm using four trajectory length schemes. We compare three fixed trajectory lengths, 200 (blue), 300 (green), and 300 with max entropy regularization (red), against Ada-NAV (ours) (yellow) in real-world environments. More details can be found in Sec. IV. The drawn trajectories are sample representations of the robot’s odometry data collected during each navigation task. Ada-NAV enhances sample efficiency of training (shown in the figure by the lower value of the orange bar plot) by adaptively varying trajectory length throughout training. Ada-NAV results in better navigation policies compared to baseline approaches and generates shorter and more successful test-time trajectories.
To improve the sample efficiency of RL for robotics navigation, common existing methods include reward-shaping [18] and learning from demonstrations [19, 20]. Reward-shaping methods use pseudo-rewards to encourage exploration in the environment [11, 21]. However, these methods require an expert to design intrinsic rewards, which can lead to performance-hindering bias. Furthermore, collecting expert-guided demonstrations for navigation tasks can be challenging and expensive to obtain [22].
@@
In this work, we present Ada-NAV, a general technique for improving the sample efficiency of RL methods for robotics navigation tasks with minimal modifications to the existing state-of-the-art methods in the literature. Our key idea is to focus on adaptively changing a salient feature in most RL methods: trajectory length. We note that in policy gradient algorithms, trajectory length is the number of interactions the robot has with the environment before each gradient update during training. For navigation, an interaction is any action the mobile robot takes and the resulting state in the environment and the reward it receives. It has been verified in the literature that short trajectory lengths can lead to inefficient learning, while long ones can be unreliable when in an unfamiliar environment [23]. Ada-NAV resolves this issue by having an adaptive trajectory length which is reliable and computationally tractable.
@@
In this work, we utilize a previously unexplored connection between a policy’s mixing time and its policy entropy to formulate Ada-NAV. To develop Ada-NAV, our novel approach is to start by connecting the trajectory length to the mixing time of the policy that generates the trajectory. The mixing time of a policy is the amount of time it takes for a robot to reach its limiting behavior under that policy. When mixing time is large, ensuring that trajectories are sufficiently long is important for obtaining accurate estimates of system behavior and algorithmic essentials, such as policy gradients [24]. When mixing time is short, on the other hand, ensuring trajectory lengths are commensurate frees up effort for performing more rapid policy updates and environmental exploration. Unfortunately, mixing time is typically difficult to estimate in complex environments [25, 26], so proxy methods for determining trajectory length are needed.
Main Contributions: Ada-NAV increases sample efficiency by controlling the amount of data required to update the policy parameter, thus decreasing the number of training samples needed to learn an optimal policy for robot navigation tasks. Our specific contributions include:
@@
  • We empirically establish a positive correlation between the policy entropy and the spectral gap of its induced Markov chain. We show that this correlation is present with various transition dynamics. This insight motivates our use of the changing policy entropy as a dynamic variable to control trajectory length.
@@
  • We take advantage of this empirical insight by utilizing a monotonic mapping between policy entropy and trajectory length to facilitate exploration in environments with sparse rewards and serve as the foundation for our adaptive trajectory length scheme, Ada-NAV.
@@
  • We present extensive empirical results using multiple RL methods, including REINFORCE [27] and Proximal Policy Optimization (PPO) [28] and Soft Actor-Critic (SAC) [29], to show that Ada-NAV improves the sample efficiency with both on-policy and off-policy methods. Our experiments in goal-reaching navigation tasks with both even and uneven terrains demonstrate that Ada-NAV outperforms constant trajectory length schemes, even those utilizing maximum entropy regularization, in terms of sample efficiency.
@@
  • We evaluate the navigation performance of Ada-NAV in both simulated and real-world outdoor environments using a Clearpath Husky robot. We observe that, for a fixed sample budget, Ada-NAV leads to an 18% increase in navigation success rate, a 20-38% decrease in the navigation path length, and a 9.32% decrease in the elevation cost compared to the policies obtained by the other methods.
@@ I-A Related Works
In this section we give a brief overview of the existing work related to RL for autonomous navigation and sparse rewards.
RL for Navigation: Motion planning and navigation has been well studied in robotics [4, 5, 6]. RL-based methods have been extensively used for robot navigations [30]. Methods such as DDPG [31], A3C [32], and DQN [33] have been used for outdoor navigation, while [34] learned a global map to find the shortest path out of a maze with RL. [35] used classical methods to create a hierarchical RL approach. For policy gradient methods, [36] demonstrated the superior performance of Soft Actor-Critic (SAC) over deep deterministic policy gradient (DDPG) in the navigation of mobile robots. In these policy gradient methods, it is well known that the gradient estimation requires lots of samples and the overall training procedure is sample inefficient.
Sparse Rewards: For model-free reinforcement learning, reward sparsity greatly affects policy search [37, 11] as a mobile robot will rarely receive any feedback to adjust its behavior. To produce more rewards which the algorithm can learn from, reward-shaping methods based on intrinsic rewards have been proposed. Methods in [38] and [21] modified the reward output to encourage exploration of the environment. [39] and [40] used imitation learning and demonstrations to overcome the challenges of sparse rewards. Curriculum learning has also been utilized to deal with sparse rewards in navigation. [41] divided the navigation task into a subgoal tree so the agent can learn from dense-to-sparse curriculum. [17] and [42] formulated a heavy-tailed policy to encourage exploration in sparse environments.
@@ II-A Reinforcement Learning and Policy Gradient Methods
We consider a Markov Decision Process (MDP) characterized by the tuple $\mathcal{M}:=(\mathcal{S},\mathcal{A},\mathbb{P},r)$, where $\mathcal{S}$ is the state space for the robot in the environment, $\mathcal{A}$ is the set of possible actions the robot can take, $\mathbb{P}$ is the transition probability kernel that determines the next state $s^{\prime}\in\mathcal{S}$ given $s\in\mathcal{S},a\in\mathcal{A}$ via $s^{\prime}\sim\mathbb{P}(\cdot|s,a)$, and $r:\mathcal{S}\times\mathcal{A}\to[0,r_{\max}]$ is the reward function. The mobile robot navigates through this environment using a (potentially stochastic) policy $\pi$ that provides a conditional probability distribution over actions $a$ given a state $s$. Given a policy $\pi$ and transition kernel $\mathbb{P}$, the transition dynamics of the Markov chain induced by $\pi$ over $\mathcal{M}$ is given by $\mathbb{P}_{\pi}(s^{\prime}|s)=\sum_{a\in\mathcal{A}}\mathbb{P}(s^{\prime}|s,a)\pi(a|s)$. The goal in the RL setting is to find a policy maximizing long-run reward over $\mathcal{M}$. To achieve this, one infinite-horizon objective typically considered in the RL literature is expected discounted reward given by
$J^{\gamma}(\pi)=\mathbb{E}_{\pi,\mu}\left[\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t})\right],$ (1)
@@
where $\gamma\in(0,1)$ is some pre-specified discount factor and $\mu$ is a given initial start state distribution over $\mathcal{S}$. In (1), the expectation is over the policy and initial state distributions. In the discounted-reward settings, the state-action value function corresponding to a given policy $\pi$ is defined by $Q^{\gamma}_{\pi}(s,a)=\mathbb{E}_{\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t})\ |\ s_{0}=s,a_{0}=a\right]$.
@@
Policy Gradient Algorithm. To handle large, potentially continuous state and action spaces, we focus on the case where policy $\pi$ is parameterized by a vector $\theta\in\mathbb{R}^{d}$, where $d$ denotes the parameter dimension, leading to the notion of a parameterized policy $\pi_{\theta}$. We define $J^{\gamma}(\theta)=J^{\gamma}(\pi_{\theta})$. Then, following policy gradient theorem [43], we note that
$\nabla J^{\gamma}(\theta)=\mathbb{E}_{\pi_{\theta},\mu}\left[Q^{\gamma}_{\pi}(s,a)\nabla\log\pi_{\theta}(a|s)\right],$ (2)
@@
which provides the gradient expression to perform gradient ascent on the objectives $J^{\gamma}(\theta)$. For the discounted setting, at time $t$ the policy gradient estimate is obtained by generating a $K\in\mathbb{N}$ length trajectory $\tau_{t}:=\{s_{t},a_{t},\ldots,s_{t+K-1}a_{t+K-1}\}$ using $\pi_{\theta_{t}}$, and then computing
$\widehat{\nabla J^{\gamma}(\theta_{t})}=\sum_{k=0}^{K-1}\gamma^{k}r(s_{t+k},a_{t+k})\nabla\log\pi_{\theta_{t}}(a_{t}|s_{t}).$ (3)
@@
This expression is then used to perform a gradient update of the form $\theta_{t+1}=\theta_{t}+\alpha\widehat{\nabla J^{\gamma}(\theta_{t})}$, for some stepsize $\alpha>0$. In REINFORCE, increasing $K$ naturally improves the quality of the estimator given in Equation (3) [44], but at the cost of an increased number of samples. In reinforcement learning for robotic navigation, simulation data to train mobile agent can be expensive to obtain or generate. Thus, reducing the number of training sample needed without degrading navigation performance is crucial in mitigating operation costs. In the next section, we will explain the quantity mixing time and its importance to trajectory length selection.
@@ II-B Mixing Time and Spectral Gap
In policy gradient algorithms an important issue is that, to obtain good estimates of the expectation (2), which is taken with respect to the stationary distribution induced over the underlying MDP by $\pi_{\theta}$, longer trajectories are often needed. The length of the trajectories is dictated by the mixing time – to be defined below – of the associated policy. Recent work on stochastic optimization [45] and policy gradient (PG) methods [24] has highlighted and clarified the connection between Markov chain mixing time and ideal trajectory length. To leverage this connection in our approach, we first present some relevant terminology and results.
We assume throughout that, for a given policy parameter $\theta$, the Markov chain induced by $\pi_{\theta}$ on the MDP $\mathcal{M}$ is irreducible and aperiodic, which is a common assumption in theoretical analyses of RL methods. First note that the ergodicity assumption guarantees that the Markov chain has a unique stationary distribution $\mu_{\theta}$. For an MDP $\mathcal{M}=(\mathcal{S},\mathcal{A},\mathbb{P},r)$ and policy $\pi_{\theta}$, the $\epsilon$-mixing time of $\pi_{\theta}$ is $\tau_{mix}^{\theta}(\epsilon):=\inf\{t:\sup_{s\in\mathcal{S}}\|\mathbb{P}_{\pi_{\theta}}^{t}(\cdot|s)-\mu_{\theta}(\cdot)\|_{TV}\leq\epsilon\}$. By convention, the mixing time is defined as $\tau_{mix}^{\theta}:=\tau_{mix}^{\theta}(1/4)$.
Though mixing times are useful in determining appropriate trajectory lengths, they are typically difficult to estimate in practice. Therefore, an alternative to estimating mixing time is estimating a lower bound. One such quantity that has been established to lower bound mixing time is the absolute spectral gap, or simply spectral gap, of the induced transition matrix $\mathbb{P}_{\pi_{\theta}}$, defined to be the difference between $1$ and the second largest eigenvalue, $\lambda^{*}_{\theta}$. As demonstrated in [46], the spectral gap and mixing time of a Markov chain bear the following relationship.
Theorem II.1
Let $\lambda_{\theta}\neq 1$ be an eigenvalue of the transition matrix $\mathbb{P}_{\pi_{\theta}}$. Then
$\tau_{mix}^{\theta}(\epsilon)\geq\left(\frac{1}{1-|\lambda_{\theta}|}-1\right)\log\left(\frac{1}{2\epsilon}\right).$ (4)
Taking $\epsilon=1/4$ and noticing that $\lambda^{*}_{\theta}$ gives the tightest lower bound, we have $\tau_{mix}^{\theta}\geq\left(1/(1-|\lambda^{*}_{\theta}|)-1\right)\log 2$. When an estimate of $\mathbb{P}_{\pi_{\theta}}$ is available, this provides a method for bounding its mixing time $\tau_{mix}^{\theta}$. Unfortunately, estimating $\mathbb{P}_{\pi_{\theta}}$ is not always practical in large and complex state and action spaces [25, 26]. Thus, a proxy for the spectral gap is crucial for navigation in complex environments. In the following section, we leverage this theorem to empirically establish a useful connection between policy entropy and spectral gap to adaptively change trajectory length with no need to estimate $\mathbb{P}_{\pi_{\theta}}$.
III Proposed Approach: Adaptive Trajectory-Based Policy Learning
Before presenting our main algorithm, Ada-NAV, we start by highlighting an interesting connection between a policy entropy and the spectral gap of its induced transition matrix.

### v7

# CCE: Sample Efficient Sparse Reward Policy Learning for Robotic Navigation via Confidence-Controlled Exploration

@@ Abstract
We introduce Confidence-Controlled Exploration (CCE), a novel exploration scheme designed to enhance the training sample efficiency of reinforcement learning (RL) algorithms for sparse reward settings such as robot navigation. Sparse rewards are common in RL and convenient to design and implement, but typically hard to deal with due to the challenges of exploration. Existing methods deploy regularization-based methods to deal with the exploration challenges. However, it is hard to characterize the balance between exploration and exploitation because regularization modifies the reward function itself, hence changing the objective we are optimizing for. In contrast to regularization-based approaches in the existing literature, our approach, CCE, is based on a novel relationship we provide between gradient estimation and policy entropy. CCE dynamically adjusts the number of samples of the gradient update used during training to control exploration. Interestingly, CCE can be applied to both existing on-policy and off-policy RL methods, which we demonstrate by empirically validating its efficacy on three popular RL methods: REINFORCE, Proximal Policy Optimization (PPO), and Soft Actor-Critic (SAC) for goal-reaching robotic navigation tasks. We demonstrate through simulated and real-world experiments that CCE outperforms conventional methods that employ constant trajectory lengths and entropy regularization when constraining the sample budget. For a fixed sample budget, CCE achieves an 18% increase in navigation success rate, a 20-38% reduction in navigation path length, and a 9.32% decrease in elevation costs. Furthermore, we showcase the versatility of CCE by integrating it with the Clearpath Husky robot, illustrating its applicability in complex outdoor environments.
@@ I Introduction
Reinforcement learning (RL) [1] is an important component of learning in robotic systems in a wide variety of complex domains, including search-and-rescue [2], disaster relief [3], and planetary exploration [4] applications. However, traditional RL methods often hinge on the availability of a dense reward structure [5], which may rarely be available in complex, unpredictable environments [6, 7, 8, 9]. Thus, reward function sparsity [10] frequently arise when applying RL methods to robotics problems, such as robotic navigation, due to their simple to implement and does not require expert knowledge to design like dense reward structures. However, this simplicity comes with a sparse feedback signal to learn from, thus leading to a limited exploration of the environment, and ultimately, to a suboptimal policy [11]. Existing literature in RL for robotics in sparse rewards has utilized methods such as entropy regularization [12], reward shaping [13, 14, 15, 16], curiosity-driven exploration [5, 17]. However, these methods are regularization-based that change the reward optimization objective which can lead to instability in learning [6].
Another major challenge of existing works comes from the sample inefficiency of exploration in sparse rewards. This issue is significant in robotic systems, due to the simulation-intensive nature of RL methods and the computational expense of realistic robotics simulation platforms [18]. Thus, RL training typically requires finding an optimal policy in a constrained sample budget. In this work, we diverge from previous exploration paradigms to respect the practicality of limited training samples. We do so by controlling exploration, rather than encouraging it based on a novel connection between the uncertainty, or confidence, of a policy and policy gradient estimation.
@@
Figure 1: Robot navigation using policies trained with the REINFORCE policy gradient algorithm using four trajectory length schemes. We compare three fixed trajectory lengths, 200 (blue), 300 (green), and 300 with max entropy regularization (red), against CCE (ours) (yellow) in real-world environments. More details can be found in Sec. IV. The drawn trajectories are sample representations of the robot’s odometry data collected. CCE enhances training sample efficiency (shown in the figure by the lower value of the orange bar plot) by adaptively varying trajectory length. CCE results in better navigation policies compared to baseline approaches and generates shorter and more successful test-time trajectories.
@@
Main Contributions: We present Confidence-Controlled Exploration (CCE), a general technique for improving the sample efficiency of RL methods for robotics navigation tasks with minimal modifications to the existing state-of-the-art methods. We introduce using the policy confidence, measured by the average conditional policy entropy over all states to control exploration. We focus on balancing exploration with the policy entropy by adaptively changing a salient feature in most RL methods: trajectory length. In policy gradient algorithms, trajectory length is the number of interactions the robot has with the environment before each gradient update during training. Our novel perspective is that exploration can be adaptively controlled by the number of samples needed to estimate the gradient. For navigation, an interaction is any action the mobile robot takes and the resulting state in the environment and the reward it receives.
@@
To design CCE, our novel approach connects the need for the policy to explore to the mixing time of the policy. The mixing time of a policy is the amount of time it takes for a robot to reach its limiting behavior under that policy. When mixing time is large, ensuring that trajectories are sufficiently long is important for obtaining accurate estimates of system behavior and algorithmic essentials, such as policy gradients [19]. When mixing time is short, on the other hand, ensuring trajectory lengths are proportional frees up effort for performing more rapid policy updates and environmental exploration. Unfortunately, mixing time is typically difficult to estimate in complex environments [20, 21], so proxy methods for determining trajectory length are needed. In this work, we utilize a previously unexplored connection between a policy’s mixing time and its policy entropy to formulate CCE. We increase the sample efficiency by adaptively tuning the trajectory length required to update the policy parameter, thus decreasing the number of samples needed to learn an optimal policy for navigation tasks. Our results include:
@@
  • We empirically establish a positive correlation between the policy entropy and the spectral gap of its induced Markov chain. We show that this correlation is present with various transition dynamics. This insight motivates our use of the changing policy entropy to control the amount of exploration during training.
@@
  • We take advantage of this empirical insight by utilizing a monotonic mapping between policy entropy and trajectory length to facilitate exploration with sparse rewards and serve as the foundation for CCE.
@@
  • We present extensive empirical results using multiple RL methods, including REINFORCE [22] and Proximal Policy Optimization (PPO) [23] and Soft Actor-Critic (SAC) [12], to show that CCE improves the sample efficiency with both on-policy and off-policy methods. Furthermore, by showing results on SAC, we show that CCE also improves sample efficiency for algorithms that employ entropy regularization. Our experiments in goal-reaching navigation tasks with both even and uneven terrains demonstrate that CCE outperforms constant trajectory length and exploration encouragement methods, with constrained samples.
@@
  • We evaluate the navigation performance of CCE in both simulated and real-world outdoor environments using a Clearpath Husky robot. We observe that, for a fixed, limited sample budget, CCE leads to an 18% increase in navigation success rate, a 20-38% decrease in the navigation path length, and a 9.32% decrease in the elevation cost compared to the policies from baselines.
@@ I-A Related Works
Sparse Rewards: For model-free reinforcement learning, reward sparsity greatly affects policy search [24, 5] as a mobile robot will rarely receive any feedback to adjust its behavior. To produce more rewards which the algorithm can learn from, reward-shaping methods based on intrinsic rewards have been proposed. Methods in [13] and [25] modified the reward output to encourage exploration of the environment. Furthermore, [26] and [27] formulated a heavy-tailed policy to encourage exploration in sparse environments. Prior work used imitation learning and demonstration [15, 28] and curriculum learning [29] to overcome the challenges of sparse rewards.
Exploration and Exploitation in RL and Navigation: To balance exploration and exploitation, prior work has used the uncertainty of a surrogate cost function estimation from the visual field of view [30] and a world model in model-based RL [31, 32]. In motion planning, [33] introduced a planner that balances exploration and exploitation using explicit information on the environmental information. By connecting exploration to policy gradient estimation, our approach, CEE, is agnostic to data modality and does not require a world model or any explicit information on the given environment. Exploration-exploitation for multi-robot systems was considered in [34, 35, 36].
RL for Navigation: Motion planning and navigation has been well studied in robotics [37, 38, 39]. RL-based methods have been extensively used for robot navigations [40]. Methods such as in [41] used DRL to find reliable paths for outdoor navigation, while [42] learned a global map to find the shortest path out of a maze with RL. In [43], classical methods were used to create a hierarchical RL approach.
@@ II-A Reinforcement Learning and Policy Gradient Methods
We consider a Markov Decision Process (MDP) characterized by the tuple $\mathcal{M}:=(\mathcal{S},\mathcal{A},\mathbb{P},r)$, where $\mathcal{S}$ is the state space for the robot in the environment, $\mathcal{A}$ is the set of possible actions the robot can take, $\mathbb{P}$ is the transition probability kernel that determines the next state $s^{\prime}\in\mathcal{S}$ given $s\in\mathcal{S},a\in\mathcal{A}$ via $s^{\prime}\sim\mathbb{P}(\cdot|s,a)$, and $r:\mathcal{S}\times\mathcal{A}\to[0,r_{\max}]$ is the reward function. The mobile robot navigates through this environment using a (potentially stochastic) policy $\pi$ that provides a conditional probability distribution over actions $a$ given a state $s$. Given a policy $\pi$ and transition kernel $\mathbb{P}$, the transition dynamics of the ergodic Markov chain induced by $\pi$ over $\mathcal{M}$ is given by $\mathbb{P}_{\pi}(s^{\prime}|s)=\sum_{a\in\mathcal{A}}\mathbb{P}(s^{\prime}|s,a)\pi(a|s)$. The goal in the RL setting is to find a policy maximizing long-run reward over $\mathcal{M}$. To achieve this, let $\mu$ be the unique stationary distribution to the ergodic Markov Chain. Thus, one infinite-horizon objective typically considered in the RL literature is the expected discounted reward given by
$J(\pi)=\mathbb{E}_{s\sim\mu,a\sim\pi}\left[\sum_{t=0}^{\infty}r(s_{t},a_{t})\right],$ (1)
@@
In (1), the expectation is over the policy and initial state distributions. In the average-reward settings, the state-action value function corresponding to a given policy $\pi$ is defined by $Q_{\pi}(s,a)=\mathbb{E}_{\pi}\left[\sum_{t=0}^{\infty}r(s_{t},a_{t})-J(\pi)\ |\ s_{0}=s,a_{0}=a\right]$.
@@
Policy Gradient Algorithm. To handle large, potentially continuous state and action spaces, we focus on the case where policy $\pi$ is parameterized by a vector $\theta\in\mathbb{R}^{d}$, where $d$ denotes the parameter dimension, leading to the notion of a parameterized policy $\pi_{\theta}$. We define $J(\theta)=J(\pi_{\theta})$. Then, following policy gradient theorem [44], we note that
$\nabla J(\theta)=\mathbb{E}_{\pi_{\theta},\mu}\left[Q_{\pi}(s,a)\nabla\log\pi_{\theta}(a|s)\right],$ (2)
@@
which provides the gradient expression to perform gradient ascent on the objectives $J(\theta)$. At update $u$ the policy gradient estimate is obtained by generating a $K\in\mathbb{N}$ length trajectory $\tau_{t}:=\{s_{0},a_{0},\ldots,s_{K-1}a_{K-1}\}$ using $\pi_{\theta_{u}}$, and then computing
$\widehat{\nabla J(\theta_{u})}=\frac{1}{K-1}\sum_{k=0}^{K-1}r(s_{k},a_{k})\nabla\log\pi_{\theta_{t}}(a_{k}|s_{k}).$ (3)
@@
This expression is then used to perform a gradient update of the form $\theta_{u+1}=\theta_{u}+\eta\widehat{\nabla J^{(}\theta_{u})}$, for some stepsize $\eta>0$. In REINFORCE, increasing $K$, and therefore the overall number of samples naturally improves the quality of the estimator given in Equation (3) [45]. However, in reinforcement learning for robotic navigation, simulation data to train mobile agent can be expensive to obtain or generate. Thus, reducing the number of training samples needed without degrading navigation performance is crucial in mitigating operation costs. In the next section, we will explain mixing time and its importance to gradient estimation.
@@ II-B Mixing Time and Spectral Gap
In policy gradient algorithms an important issue is that, to obtain good estimates of the expectation in (2), which is taken with respect to the stationary distribution, $\mu_{\theta}$, induced over the underlying MDP by $\pi_{\theta}$, longer trajectories are often needed. The length of the trajectories is dictated by the mixing time – to be defined below – of the associated policy. Recent work on stochastic optimization [46] and policy gradient (PG) methods [19] has highlighted and clarified the connection between Markov chain mixing time and ideal trajectory length, which we will later use to connect confidence to exploration.
We assume throughout that, for a given policy parameter $\theta$, the Markov chain induced by $\pi_{\theta}$ on the MDP $\mathcal{M}$ is ergodic, which is a common assumption in theoretical analyses of RL methods [19, 47, 46, 48]. First note that the ergodicity assumption guarantees that the Markov chain has a unique stationary distribution $\mu_{\theta}$. For an MDP $\mathcal{M}=(\mathcal{S},\mathcal{A},\mathbb{P},r)$ and policy $\pi_{\theta}$, the $\epsilon$-mixing time of $\pi_{\theta}$ is $\tau_{mix}^{\theta}(\epsilon):=\inf\{t:\sup_{s\in\mathcal{S}}\|\mathbb{P}_{\pi_{\theta}}^{t}(\cdot|s)-\mu_{\theta}(\cdot)\|_{TV}\leq\epsilon\}$. By convention, the mixing time is defined as $\tau_{mix}^{\theta}:=\tau_{mix}^{\theta}(1/4)$ [47].
Though mixing times are useful in determining trajectory lengths, they are difficult to estimate in practice. Therefore, an alternative is estimating a lower bound. As described in [49], one such lower bound of mixing time is the absolute spectral gap, or simply spectral gap, of the induced transition matrix $\mathbb{P}_{\pi_{\theta}}$, defined to be the difference between $1$ and the second largest eigenvalue, $\lambda^{2}_{\theta}$:
When an estimate of $\mathbb{P}_{\pi_{\theta}}$ is available, this provides a method for bounding its mixing time. Unfortunately, estimating $\mathbb{P}_{\pi_{\theta}}$ is impractical in large and complex state and action spaces [20, 21]. Thus, a proxy for the spectral gap is crucial for navigation in complex environments. In the following section, we leverage this theorem to empirically establish a useful connection between policy entropy and spectral gap to control the exploration with no need to estimate $\mathbb{P}_{\pi_{\theta}}$.
III Proposed Approach: Controlled Exploration-based Policy Learning